OpenStudy (dmezzullo):

What does it mean for a number to be the "solution" to an equation? Give an example of a variable equation and its solution. Explain in complete sentences how you know that number is the solution.

OpenStudy (dmezzullo):

@mikala1

OpenStudy (anonymous):

A solution of an equation is any entity that can be substituted for the variable or variables such that the the equation is a true statement. For example, consider the equation It should be clear that if you replace the with , then you have a true statement because 5 plus 3 does indeed equal 8. But if you replace with any other number, then you have a false statement. Let's see what happens if we use . 2 plus 3 is 5 which is most assuredly not equal to 8. And the same thing would happen for any other number besides 5 substituted for .

OpenStudy (anonymous):

the example is x pluse 3 is 8

OpenStudy (anonymous):

@dmezzullo a no. is a solution means it satisfies the equation example-----2x+3=15 means a no. which satisfy it 2x=15-3 2x=12 x=6 when we put 6 in original equation it satisfies it LHS=2(6)+3=12+3=15=RHS :)